# Position-building in competition with real-world constraints

### v1

@@ 1.1 Definitions and terminology
In Section 5 we introduce an algorithm for computing two-trader equilibria that in effect replicates this dynamic adjustment process (see 1). The algorithm allows one to understand the entire path leading to equilibrium and helps to explain certain features of equilibria such as in certain cases the cost of trading in equilibrium are higher than non-equilibrium, which can be seen as a direct consequence of the non-cooperative nature of the games.
@@ 1.2 Mathematical overview of trading strategies
$\text{Cost}(a,b_{\lambda})=\int_{0}^{1}(\dot{a}+\lambda\dot{b}_{\lambda})+\kappa(a+\lambda b_{\lambda})\dot{a}\,\texttt{dt}$ (1.2)
@@
$\underset{\underset{a(0)=0}{a:[0,1]\to\mathbf{R}}}{\operatorname{\mathbf{minimize}}}\,\,\int_{0}^{1}(\dot{a}+\lambda\dot{b}_{\lambda})\dot{a}+\kappa(a+\lambda b_{\lambda})\dot{a}\,\texttt{dt}$ (1.3)
@@ 2.1 Approximating the cost function with Fourier series
$\begin{split}&\text{Cost}_{\mathcal{F}}(a,b_{\lambda})=\\ &\quad\frac{1}{2}(2+\kappa)(1+\lambda)+\frac{\pi^{2}}{2}\sum_{n=1}^{N}(a_{n}^{2}+\lambda a_{n}b_{n})n^{2}+\frac{2\kappa\lambda}{\pi}\sum_{\begin{subarray}{c}n=1\\ n\,\text{odd}\end{subarray}}^{N}\frac{b_{n}-a_{n}}{n}+2\kappa\!\!\!\!\sum_{\begin{subarray}{c}n,m=1\\ n+m\,\text{odd}\end{subarray}}^{N}\frac{(a_{m}+\lambda b_{m})a_{n}nm}{m^{2}-n^{2}}\end{split}$ (2.1)
@@
$\begin{split}&\text{Cost}_{\mathcal{F},b_{\lambda}}(a)=\\ &\quad\lambda\left\{\frac{1}{2}(2+\kappa)(1+\lambda)+\frac{\pi^{2}}{2}\sum_{n=1}^{N}\left(\lambda b_{n}^{2}+a_{n}b_{n}\right)n^{2}+\frac{2\kappa}{\pi}\sum_{\begin{subarray}{c}n=1\\ n\,\text{odd}\end{subarray}}^{N}\frac{a_{n}-b_{n}}{n}+2\kappa\!\!\!\!\sum_{\begin{subarray}{c}n,m=1\\ n+m\,\text{odd}\end{subarray}}^{N}\frac{(a_{m}+\lambda b_{m})b_{n}nm}{m^{2}-n^{2}}\right\}\end{split}$ (2.2)
@@ 2.2 Properties of the approximate cost function
The approximate cost function $\text{Cost}_{\mathcal{F},a}(b_{\lambda})$ in equation (LABEL:eq:fourier-cost-fn) is quadratic in the Fourier coefficients of $a_{\mathcal{F}}$, and similarly $\text{Cost}_{\mathcal{F},b_{\lambda}}(a)$ is quadratic in the coefficients of $b_{\mathcal{F}}$. Also, an immediate consequence of LABEL:eq:fourier-cost-fn and LABEL:eq:fourier-cost-fn-ba-final is that the Hessian matrix of both are diagonal and positive definitive, which will imply that the quadratic programs we focus on when adding constraints will be convex. Both cost functions depend implicitly on the number of Fourier coefficients used to approximate $a_{\mathcal{F}}$ and $b_{\mathcal{F}}$ and we can show that as $N\to\infty$ the approximate cost functions converge to the actual cost functions. We state this now and sketch the proof.
@@ 3.1 How to constrain best-response and equilibrium strategies
Approximating an interval with a finite set of points: Our aim below is to constrain a function $f$ on an interval $I_{ab}=[a,b]$ well enough that the size of the set of points where a constraint such as $f(t)<g(t)$ for all $t\in I_{ab}$ fails to hold is sufficiently small. Thus we want a set $\mathcal{T}=\{t_{1},\dots,t_{K}\}\subset\mathcal{T}$ such that
@@ 3.2 Overbuying constraint
Figure 1: An example of a over-buying strategy (blue line) exceeds the target quantity of one and therefore violates the over-buying constraint.
@@ 3.6 No-sell constraints
$\begin{array}[]{ll}\operatorname{\mathbf{minimize}}&\quad\text{Cost}_{\mathcal{F},a}(b_{\lambda})\\[5.38193pt] \text{subject to}&\quad t+\sum_{n=1}^{N}a_{n}n\pi\cos(n\pi t)\geq 0t\in\mathcal{T}\end{array}$ (3.8)
@@ 4 Numerical examples of strategies with constraints
In the sections that follow we provide numerical examples of best-response and equilibrium strategies with constraints. We run through a collection of strategies that a single adversary might trade. [^3] The code for all of these examples can be found online at https://github.com/vragulin/Optimal_Exec_In_Comp, commit 8dbb242. To explain this start by noting that every adversary depends on a scaling parameter $\lambda$, those strategies that take into account the market impact coefficient $\kappa$ will depend on $\kappa$ and those arise when, for instance, the strategy is itself a best-response strategy. Finally many strategies are themselves members of parametric families, for example risk-averse strategies depend on a risk aversion parameter, $\sigma$ (see Section 4.3).
@@ 4.1 A taxonomy of trading strategies
    Best-response: are optimal response strategies when trading in competition with some other strategy. The operative question when investigating a best-response strategy is what is it the best response to? These strategies are require knowledge of both what the strategy that is being responded to is, and the market impact coefficient $\kappa$. Therefore, as a starting place each passive strategy has a best-response strategy.
@@
    • Risk-neutral: Best-response strategies to risk-averse strategies and are discussed in Section 4.2.
@@ 4.2 Best-response to a risk-neutral adversary
$a(t;{\kappa,\!\lambda\!})=\left(1-\frac{\lambda\kappa}{4}\right)t^{2}+\frac{\lambda\kappa}{4}t$ (4.1)
@@
Figure 4: Best-response strategies $a$ to a risk-neutral adversary trading $b_{\lambda}$ with constrained to 500% and 200% of the target quantity. The plots show unconstrained (solid red) and overbuy constrained (dashed red) strategies, limiting quantities held to less than $1+\rho$ for various values of $\rho$.
@@ 4.3 Best response to a risk-averse adversary
$a(t;\sigma)=\dfrac{\lambda}{2}(q_{\sigma}(0)-q_{\sigma}(t))+\big{(}1+\dfrac{\lambda}{2}(q_{\sigma}(1)-q_{\sigma}(0))\big{)}\,t$ (4.2)
@@
Figure 5: Best-response strategies to a risk-averse competitor $b_{\lambda}(t)$ with risk-aversion parameter $\sigma$ as in Eq. 4.2 for $\sigma=3$. The results are the strategy $a$ both unconstrained (solid red) and with the constrained for various values of $\rho$.
@@ 4.4 Best response to an eager adversary
Figure 6: Best-response strategies to an eager competitor $b_{\lambda}(t)$ (in blue) with eagerness parameter $\sigma$=3 as in Eq. 4.5 for all plots. The best-response strategies are in solid and dashed red, where the solid red lines are unconstrained and the dashed red are various constraints on Short-selling, see Section 3.5. In these plots the trader must complete the fraction $c$ of the strategy by a time $t$ and then remain at or below the target quantity. The grey boxes show the "end-strategy region" for each strategy in each plot.
@@
Figure 7: Best-response strategies to an eager competitor $b_{\lambda}(t)$ with eagerness parameter $\sigma$ as in Eq. 4.5 for $\sigma=3$. The results are strategy $a(t)$ with end-strategy constraints (see Section 3.4).
@@ 4.5 Two-trader equilibrium closed-form solutions
If $A$ thinks that $B$ will trade the two-trader equilibrium strategy Eq. 4.6b for a given $\kappa$ and $\lambda$ then $A$’s best-response is to trade $a_{\texttt{eq}}(t;{\kappa,\!\lambda\!})$. If, however, $A$ needs to add constraints, e.g. an constraint, then while $A$ can add a constraint and solve for the best-response to $b(t;\kappa,\lambda b_{\lambda})$, this will not be an equilibrium strategy, because $B$’s strategy is not the best response to $A$’s constrained best-response to $B$. Nevertheless, Figure 8 depicts these best responses.
Figure 8: Best-response strategies for trader $A$ assuming that trader $B$ trades the two-trader equilibrium strategy $b(t;{\kappa,\!\lambda\!})$, while $A$ complies with a channel constraint set at various levels (see Section 3.3) requiring that $a(t)$ is always between the risk-neutral lower bound and the specified eager benchmark upper bound (4.4).
@@ 5 Computing two-trader equilibrium via iterative quadratic programming
$\begin{array}[]{cl}\underset{(a_{1}\cdots a_{N})}{\operatorname{\mathbf{minimize}}}&\quad\text{Cost}_{\mathcal{F}}(a;b_{\lambda}^{(k)})\\[8.61108pt] \text{subject to}&\quad{\mathfrak{C}}(a;c)\end{array}$ (5.5) $\begin{array}[]{cl}\underset{(b_{1}\cdots b_{N})}{\operatorname{\mathbf{minimize}}}&\quad\text{Cost}_{\mathcal{F}}(b_{\lambda},a^{(k)})\\[8.61108pt] \text{subject to}&\quad{\mathfrak{C}}(b;d)\end{array}$ (5.6)
@@
Step 1(i):Minimize Cost(a.bλ(0)) s.t. (a;c)Produces: a(1)Step 1(ii):Minimize Cost(bλ,a(1)) s.t. (b;d) Produces: bλ(1) $\begin{array}[]{lll}\textbf{Step 1(i):}&\quad\text{Minimize $\text{Cost}_{\mathcal{F}}(a.b_{\lambda}^{(0)})$ s.t. ${\mathfrak{C}}(a;c)$}&\quad\textbf{{Produces}: $a^{(1)}$}\\ \textbf{Step 1(ii):}&\quad\text{Minimize $\text{Cost}_{\mathcal{F}}(b_{\lambda},a^{(1)}$) s.t. ${\mathfrak{C}}(b;d)$ }&\quad\textbf{Produces: $b_{\lambda}^{(1)}$ }\end{array}$ (5.7)
@@
5:     $\quad a^{(k)}\leftarrow\gamma a+(1-\gamma)a^{(k\shortminus 1)}$
@@
Note that in each successive step in 1 we do not optimize using the prior steps output, but rather we step toward the function along the line that interpolates in Fourier coefficient space between the prior step’s strategy and the best-response strategy. We find that in practice this leads to significantly better convergence. In what follows we overload the term best-response to include these interpolations. See Figure 15 for an illustration of the impact of $\gamma$ on the equilibrium path.
@@ 6.1 Computing equilibria and the dynamic path to equilibrium
We begin by studying two-trader equilibria with no constraints to achieve to purposes. With no constraints we can compare the optimization results to the closed form solutions in (4.6a) and (4.6b); and we can study the path to convergence build up from each step in Algorithm 1. This provides insight into the nature of the dynamic process of adjustments that leads to convergence.
@@
Plots: Figures 9 to 11 provide three illustrations of solving the unconstrained two-trader equilibrium using Algorithm 1. Each optimization starts with the initial guess $a^{(0)}=t$, the risk-neutral strategy and the first iteration is $b_{\lambda}^{(1)}$ the best-response to this strategy. In all of these plots $\gamma=0.2$ (see 1). The upper left-hand plot of each shows the trading strategies for $a$ and $b_{\lambda}$ on top of one another; the upper right-hand plot shows the difference between the closed-form and approximate solutions for both the initial iteration and the final iteration of the optimization. For example, in Figure 9 ($\lambda=5$ and $\kappa=1$) the right hand plot shows in solid red the plot of $b^{(1)}(t)-b_{\texttt{eq}}(t)$ which shows that at approximately $t=0.5$ the equilibrium strategy strategy holds approximate 0.35 units more than the initial guess (the risk-neutral strategy). The initial guess for $a$ is a lot closer in absolute terms (dark blue) but with $\lambda=5$ this scale difference makes sense.

### v2

@@ 1.1 Definitions and terminology
In the case of position building in competition this is exactly what happens when we introduce computational methods for finding two-player equilibria. In Section 5 we introduce an algorithm for computing two-trader equilibria that in effect replicates this dynamic adjustment process (see 1). The algorithm allows one to understand the entire path leading to equilibrium and helps to explain certain features of equilibria such as in certain cases the cost of trading in equilibrium are higher than non-equilibrium, which can be seen as a direct consequence of the non-cooperative nature of the games.
@@ 1.2 Mathematical overview of trading strategies
$\text{Cost}(a,b_{\lambda})=\int_{0}^{1}(\dot{a}+\lambda\dot{b}_{\lambda})\dot{a}+\kappa(a+\lambda b_{\lambda})\dot{a}\,\texttt{dt}$ (1.2)
@@
$\underset{\underset{a(0)=0,a(1)=1}{a:[0,1]\to\mathbf{R}}}{\operatorname{\mathbf{minimize}}}\,\,\int_{0}^{1}(\dot{a}+\lambda\dot{b}_{\lambda})\dot{a}+\kappa(a+\lambda b_{\lambda})\dot{a}\,\texttt{dt}$ (1.3)
@@
1.3 Convex combinations of trading strategies and the cost function [^3] This section is new to Version 2. The prior version stated Algorithm 1 without any explanation of why it works.
Consider $0\leq\gamma\leq 1$ and two strategies $b_{\lambda,1}$ and $b_{\lambda,2}$ we can form the convex combination
$b_{\lambda,\gamma}=\gamma b_{\lambda,1}+(1\shortminus\gamma)b_{\lambda,2}$ (1.11)
and then for a strategy $a$ we have the associated cost functions $\text{Cost}(a,\gamma b_{\lambda,1})$, $\text{Cost}(a,(1\shortminus\gamma)b_{\lambda,2})$ and $\text{Cost}(a,b_{\lambda,\gamma})$. It is immediate consequence of the linearity of integration and differentiation and they are related as follows:
$\text{Cost}(a,b_{\lambda,\gamma})=\text{Cost}(a,\gamma b_{\lambda,1})+\text{Cost}(a,(1\shortminus\gamma)b_{\lambda,2})$ (1.12)
The meaning of Eq. 1.12 is as follows. The left-hand side is the cost of $a$ trading in competition with $b_{\lambda,\gamma}$, while the right-hand side is the sum of the cost of $a$ trading in competition with $\gamma b_{\lambda,1}$ and $a$ trading in competition with $(1\shortminus\gamma)b_{\lambda,2}$. Now suppose that $a^{*}$ is optimal versus $b_{\lambda,\gamma}$, that is, that it minimizes $\text{Cost}(a,b_{\lambda,\gamma})$.
@@ 2.1 Approximating the cost function with Fourier series
$\text{Cost}_{\mathcal{F}}(a,b_{\lambda})=\frac{1}{2}(2+\kappa)(1+\lambda)+\frac{\pi^{2}}{2}\sum_{n=1}^{N}(a_{n}^{2}+\lambda a_{n}b_{n})n^{2}+\frac{2\kappa\lambda}{\pi}\sum_{\begin{subarray}{c}n=1\\ n\,\text{odd}\end{subarray}}^{N}\frac{b_{n}-a_{n}}{n}+2\kappa\!\!\!\!\sum_{\begin{subarray}{c}n,m=1\\ n+m\,\text{odd}\end{subarray}}^{N}\frac{\lambda b_{m}a_{n}nm}{m^{2}-n^{2}}$ (2.1)
@@
$\text{Cost}_{\mathcal{F},b_{\lambda}}(a)=\lambda\left\{\frac{1}{2}(2+\kappa)(1+\lambda)\!+\!\frac{\pi^{2}}{2}\sum_{n=1}^{N}\left(\lambda b_{n}^{2}\!+\!a_{n}b_{n}\right)n^{2}+\frac{2\kappa}{\pi}\sum_{\begin{subarray}{c}n=1\\ n\,\text{odd}\end{subarray}}^{N}\frac{a_{n}-b_{n}}{n}\!+\!2\kappa\!\!\!\!\sum_{\begin{subarray}{c}n,m=1\\ n+m\,\text{odd}\end{subarray}}^{N}\frac{a_{m}b_{n}nm}{m^{2}-n^{2}}\right\}$ (2.2)
@@ 2.2 Properties of the approximate cost function
The approximate cost function $\text{Cost}_{\mathcal{F},a}(b_{\lambda})$ in equation (2.1) is quadratic in the Fourier coefficients of $a_{\mathcal{F}}$, and similarly $\text{Cost}_{\mathcal{F},b_{\lambda}}(a)$ is quadratic in the coefficients of $b_{\mathcal{F}}$. Also, an immediate consequence of Eq. 2.1 and Eq. 2.2 is that the Hessian matrix of both are diagonal and positive definitive, which will imply that the quadratic programs we focus on when adding constraints will be convex. Both cost functions depend implicitly on the number of Fourier coefficients used to approximate $a_{\mathcal{F}}$ and $b_{\mathcal{F}}$ and we can show that as $N\to\infty$ the approximate cost functions converge to the actual cost functions. We state this now and sketch the proof.
@@
2.3 Some useful facts about Fourier series of trading strategies [^5] This section is new to Version 2.
There are a number of basic properties of the Fourier series of trading strategies that we note here.
Convex combinations of trading strategies. Given unit strategies $a_{1},a_{2}$ and $0\leq\gamma\leq 1$, $a_{12}=\gamma a_{1}+(1\shortminus\gamma)a_{2}$ is a convex combination of the strategies. Since $a_{12}(0)=0$ and $a_{12}(1)=1$, it has a Fourier series:
$a_{12,\mathcal{F}}=t+\sum_{n=1}^{\infty}a_{12,n}\sin(n\pi t)$ (2.4)
We can also form the Fourier series of the function $\gamma a_{1}+(1\shortminus\gamma)a_{2}$:
$\gamma a_{1}+(1\shortminus\gamma)a_{2}=t+\sum_{n=1}^{\infty}\left(\gamma a_{1,n}+(1-\gamma)a_{2,n}\right)\sin(n\pi t)$ (2.5)
and it is immediately clear that the series in Eq. 2.5 the expression in Eq. 2.4 and Eq. 2.5 must be equal which immediately implies:
$\gamma a_{1,n}+(1\shortminus\gamma)a_{2,n}=a_{12,n},\qquad\text{for all $n$}$ (2.6)
We will use this fact in Section 4.5.
@@ 3.1 How to constrain best-response and equilibrium strategies
Approximating an interval with a finite set of points: Our aim below is to constrain a function $f$ on an interval $I_{ab}=[a,b]$ well enough that the size of the set of points where a constraint such as $f(t)<g(t)$ for all $t\in I_{ab}$ fails to hold is sufficiently small. Thus we want a set $\mathcal{T}=\{t_{1},\dots,t_{K}\}\subset[a,b]$ such that
@@ 3.2 Overbuying constraint
Figure 1: An example of a over-buying strategy (blue line) exceeds the target quantity of one and therefore violates the overbuy constraint.
@@ 3.6 No-sell constraints
$\begin{array}[]{ll}\operatorname{\mathbf{minimize}}&\quad\text{Cost}_{\mathcal{F},a}(b_{\lambda})\\[5.38193pt] \text{subject to}&\quad t+\sum_{n=1}^{N}a_{n}n\pi\cos(n\pi t)\geq 0,\,t\in\mathcal{T}\end{array}$ (3.8)
@@ 4 Numerical examples of strategies with constraints
In the sections that follow we provide numerical examples of best-response and equilibrium strategies with constraints. We run through a collection of strategies that a single adversary might trade. [^7] The code for all of these examples can be found online at https://github.com/vragulin/Optimal_Exec_In_Comp, commit 7bc0c74. To explain this start by noting that every adversary depends on a scaling parameter $\lambda$, those strategies that take into account the market impact coefficient $\kappa$ will depend on $\kappa$ and those arise when, for instance, the strategy is itself a best-response strategy. Finally many strategies are themselves members of parametric families, for example risk-averse strategies depend on a risk aversion parameter, $\sigma$ (see Section 4.3).
@@ 4.1 A taxonomy of trading strategies
    Best-response: are optimal response strategies when trading in competition with some other strategy. The operative question when investigating a best-response strategy is what is it the best response to? These strategies require knowledge of both what the strategy that is being responded to is, and the market impact coefficient $\kappa$. Therefore, as a starting place each passive strategy has a best-response strategy.
@@
    • Risk-neutral: Best-response strategies to risk-neutral strategies and are discussed in Section 4.2.
@@ 4.2 Best-response to a risk-neutral adversary
$a(t)=\left(1+\frac{\lambda\kappa}{4}\right)t-\frac{\lambda\kappa}{4}t^{2}$ (4.1)
@@
Figure 4: Best-response strategies $a$ to a risk-neutral adversary trading $b_{\lambda}$ with an overbuying constraint. The plots depict two different overbuying buying constraints (500% and 200% of the target quantity) in dashed red and the unconstrained in solid red.
@@ 4.3 Best response to a risk-averse adversary
$a(t)=\dfrac{\lambda}{2}(q_{\sigma}(0)-q_{\sigma}(t))+\big{(}1+\dfrac{\lambda}{2}(q_{\sigma}(1)-q_{\sigma}(0))\big{)}\,t$ (4.2)
@@
Figure 5: Best-response strategies $a$ to a risk-averse competitor trading $b_{\lambda}$ with risk-aversion parameter $\sigma$ as in Eq. 4.2 for $\sigma=3$. The results are shown for both unconstrained (solid red) and with the constrained for various values of $\rho$.
@@ 4.4 Best response to an eager adversary
Figure 6: Best-response strategies to an eager competitor $b_{\lambda}(t)$ (in blue) with eagerness parameter $\sigma$=3 as in Eq. 4.5 for all plots. The best-response strategies are in solid and dashed red, where the solid red lines are unconstrained and the dashed red are various constraints on Short-Selling, see Section 3.5. In these plots the trader is not allowed to establish a position below $c$ at any time, i.e. $a(t)\geq c,\,\text{for all }t\in[0,1]$.
@@
Figure 7: Best-response strategies to a passive eager competitor trading $b_{\lambda}$ with eagerness $\sigma$ (see Eq. 4.5) for $\sigma=3$ and $\kappa=0.5$ computed with a variety of end-strategy constraints depicted by the grey shaded boxes (see Section 3.4).
@@ 4.5 Two-trader equilibrium closed-form solutions
If $A$ thinks that $B$ will trade the two-trader equilibrium strategy Eq. 4.6b for a given $\kappa$ and $\lambda$ then $A$’s best-response is to trade $a_{\texttt{eq}}(t)$. If, however, $A$ requires constraints on its strategy then while $A$ can add a constraint and solve for the best-response to $b_{\lambda}$, this will not be an equilibrium strategy, because $B$’s strategy is not the best response to $A$’s constrained best-response to $B$. Nevertheless, Figure 8 depicts these best responses.
Figure 8: Best-response strategies to the two-trader equilibrium strategy $b_{\lambda}$ for various values of $\lambda$; the best-response $a$ is constrained by the channel constraint set at various levels as depicted by the shaded grey areas (see Section 3.3). The channel constraints require that $a$ is always between the risk-neutral lower bound and the specified eager benchmark upper bound (4.4).
@@ 5 Computing two-trader equilibrium via iterative quadratic programming
$\begin{array}[]{cl}\underset{(a_{1}\cdots a_{N})}{\operatorname{\mathbf{minimize}}}&\quad\text{Cost}_{\mathcal{F}}(a,b_{\lambda}^{(k)})\\[8.61108pt] \text{subject to}&\quad{\mathfrak{C}}(a;c)\end{array}$ (5.5) $\begin{array}[]{cl}\underset{(b_{1}\cdots b_{N})}{\operatorname{\mathbf{minimize}}}&\quad\text{Cost}_{\mathcal{F}}(b_{\lambda},a^{(k)})\\[8.61108pt] \text{subject to}&\quad{\mathfrak{C}}(b;d)\end{array}$ (5.6)
@@
Step 1(i):Minimize Cost(a,bλ(0)) s.t. (a;c)Produces: a(1)Step 1(ii):Minimize Cost(bλ,a(1)) s.t. (b;d) Produces: bλ(1) $\begin{array}[]{lll}\textbf{Step 1(i):}&\quad\text{Minimize $\text{Cost}_{\mathcal{F}}(a,b_{\lambda}^{(0)})$ s.t. ${\mathfrak{C}}(a;c)$}&\quad\textbf{{Produces}: $a^{(1)}$}\\ \textbf{Step 1(ii):}&\quad\text{Minimize $\text{Cost}_{\mathcal{F}}(b_{\lambda},a^{(1)}$) s.t. ${\mathfrak{C}}(b;d)$ }&\quad\textbf{Produces: $b_{\lambda}^{(1)}$ }\end{array}$ (5.7)
@@
5:     $\quad a^{(k)}\leftarrow\gamma a+(1\shortminus\gamma)a^{(k\shortminus 1)}$
@@
Note that in each successive step in 1 optimize by minimizing the cost function trading not versus the prior best-response strategy but rather versus a convex combination of the prior best-response and current best-response’s Fourier coefficients. As we saw in Section 2.3, the convex combination of the strategy’s Fourier coefficients is an approximation of the same convex combination of strategies. We find that in practice this leads to significantly better convergence. In what follows we overload the term best-response to include these interpolations. See Figure 15 for an illustration of the impact of $\gamma$ on the equilibrium path.
@@ 6.1 Computing equilibria and the dynamic path to equilibrium
We begin by studying two-trader equilibria with no constraints to achieve two purposes. With no constraints we can compare the optimization results to the closed form solutions in (4.6a) and (4.6b); and we can study the path to convergence build up from each step in Algorithm 1. This provides insight into the nature of the dynamic process of adjustments that leads to convergence.
@@
Plots: Figures 9 to 11 provide three illustrations of solving the unconstrained two-trader equilibrium using Algorithm 1. Each optimization starts with the initial guess $a^{(0)}=t$, the risk-neutral strategy and the first iteration is $b_{\lambda}^{(1)}$ the best-response to this strategy. Unless otherwise stated, these plots $\gamma=0.8$ (see 1). We use a lower value when convergence is an issue. The upper left-hand plot of each shows the trading strategies for $a$ and $b_{\lambda}$ on top of one another; the upper right-hand plot shows the difference between the closed-form and approximate solutions for both the initial iteration and the final iteration of the optimization. For example, in Figure 9 ($\lambda=5$ and $\kappa=1$) the right hand plot shows in solid red the plot of $b^{(1)}(t)-b_{\texttt{eq}}(t)$ which shows that at approximately $t=0.5$ the equilibrium strategy strategy holds approximate 0.35 units more than the initial guess (the risk-neutral strategy). The initial guess for $a$ is a lot closer in absolute terms (dark blue) but with $\lambda=5$ this scale difference makes sense.